the frequency of a rare, autosomal recessive metabolic disease in the U.S. population is 1 in 50,000. assuming h-w equilibrium, what percentage of our population should be carriers (heterozygous) for this disease?
Answer #1

Given,

The frequency of rare autosomal recessive diseases = q2 = 1/ 50,000

Therefore, q= 0.00447

Now by H--W Equilibrium

p+q=1

by using above equation we can calculate p.

p=1-0.00447

=0.9955

As we know that heterozygous frequency of population=2pq

=2*0.9955*0.00447

=0.00889977

Therefore the percentage of heterozygous population=0.889977%